# Testing identifying assumptions in Tobit Models

### v1

@@ Abstract.
This paper develops sharp testable implications for Tobit and IV-Tobit models’ identifying assumptions: linear index specification, (joint) normality of latent errors, and treatment (instrument) exogeneity and relevance. The new sharp testable equalities can detect all possible observable violations of the identifying conditions. We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds. Simulation results suggests proper size for large samples and that the test is powerful to detect large violation of the exogeneity assumption and violations in the error structure. Finally, we review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions.
@@ 1. Introduction
Since the seminal work of Tobin (1958), Tobit models have earned attention in economics, business and social sciences. [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations. Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative. This approach is related to a broader class of models known as censored or truncated regression models depending on whether observations outside a specified range are lost or censored. When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE). Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and (iii), in the case of an instrumental variable approach to endogeneity, the relevance condition for the instrument. While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
@@
In this paper, we provide a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature. The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model. Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability. Following recent literature, we convert the equalities into conditional moment inequalities, and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013). The third contribution is to review and propose alternative approaches that can be used when the model is rejected. We explore an alternative path to partially identify the parameter of interest by assuming monotonicity of selection into treatment. Finally, we provide empirical examples illustrating the methodology and its practical relevance. More generally, the current paper contributes to the growing literature on testing identifying assumptions of econometric models.
@@
We focus on two main models: (i) the “classic Tobit” model in which the main variable of interest is assumed to be exogenous and (ii) the IV-Tobit model. In both cases, the proposed test considers all observable violations of the model structure, such as linear index, normality of the latent errors, independence of the treatment, and homoskedasticity; with the addition of the validity of the instrument for the IV-Tobit. Additional results for variants of the Tobit family of models (e.g., type 2 Tobit) are presented in the Appendix.
@@ 1.1. Previous Literature
There is a vast literature related to testing the validity of Tobit models and their assumptions, to which we contribute. Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid. Nelson (1981) constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates. Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model. As stated by Bera, Jarque, and Lee (1984), Nelson (1981) test is equivalent to the Lagrange multiplier test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984). Bera, Jarque, and Lee (1984) proposes a Lagrange Multiplier test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained. Newey (1987) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test. Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the Lagrange multiplier (score) test for a common null hypothesis. Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
@@
Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model. With a similar intuition to our framework, their test writes down conditional moment restrictions which should have zero conditional expected values under the null. The intuition behind the conditional moments test developed by Drukker (2002) is that since the model was estimated by maximum likelihood, the assumed data-generating process specifies all moments of disturbances conditional on the covariates to be the ones of a normal distribution. Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution. Our procedure, while intuitively similar detects all possible violations of the model, not only those evident from deviations in the third and fourth moments. Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables.
Most of these approaches focus on testing a particular subset of assumptions, or a specific class of alternatives to consider. We consider the full set of assumptions with a general alternative hypothesis, which is not a specific subset of models.
@@
The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data. This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models. Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, etc (Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023). Our testing procedures are connected to kedagni2020generalized which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure which are inherent to Tobit models.
@@
The remainder of the paper is organized as follows. Section 2 presents the model and identifying assumptions. Section 3 derives the sharp testable implications. Section 4 discusses heuristically the identification of the model’s parameters. Section 5 outlines the testing procedure. Section 6 include simulation evidence about the test’s size and power. Section 7 discusses how to relax the assumptions in case of rejection. Section 8 and 9 provide empirical illustrations. Finally, Section 10 concludes. Additional results and details are collected in Appendices A-E.
@@ 2. Models
This section describes two popular “two-part” models for truncated data for which we derive testable implications in Section 3: the classic Tobit, and the instrumental variable Tobit. [^2] We omit conditioning on other exogenous covariates $X$ for simplicity of exposition.
@@ 2.1. Classic Tobit model
Where $U$ is an unobservable (latent) error. In addition to the model structure in Equation 2.4, the classic Tobit also considers restrictions on the distribution of the error term and its relationship to the treatment.
@@
Assumption 1 states that the treatment of interest is independent of the model’s unobservables. Assumption 2 imposes that the latent error has a standard normal distribution. [^4] The error could take $N(0,\sigma^{2})$ but after scale normalizations of the coefficients of $Y^{*}$ we can model it as $N(0,1)$.
Despite its simplicity, the classic Tobit model can accommodate treatment that is continuous or discrete, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
@@ 2.2. IV Tobit model
The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect. Let the IV Tobit model be: [^5] Results for the IV-Tobit model are presented for continuous $D$ but can accommodate a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$. In that case, the parameters are identified only up to scale.
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.10)
@@
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.16)
Where $U,V$ are the latent structural error terms. To that model structure, we add restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable.
@@
Assume that: $Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
@@
Assumption 4 characterises the distribution of the latent vector of error terms as bivariate normal. Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. [^6] Note that if $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$. Then the reduced form system can be scale normalized. The latter characterisation aids the development of the sharp testable implications.
@@ 3. Sharp testable equalities
This section presents the sharp testable equalities for the models described in Section 2. Deviations from these equalities imply violations of the “null hypothesis” of the Tobit model (linear latent index, independence, instrument validity, and normality). The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 4.
@@ 3.1. Classic Tobit model
Starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ are observed. For any value of the treatment variables $d\in\mathcal{D}$:
@@
Turning to the accumulation point, the observed event of $Y=0$,
$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=1-\Phi(\alpha_{0}+\alpha_{1}d).$ (3.3)
@@
In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$. For the proof see Appendix A.
@@
The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable. However, the model is non-verifiable. To prove this, we show that for any joint distribution on the data $(Y,D)$ that satisfies equalities 3.1-3.3, there exists a vector of $(\bar{Y},D,\bar{U})$ that induces the observed distribution on the data $(Y,D)$, but the vector $\bar{U}$ is not normally distributed. See Appendix E for the proof.
@@
The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (barros2018generalized), different thresholds (carson2007Tobit), dynamic Tobit models, or including individual specific effects (wooldridge2005simple; honore2000estimation; honore1993orthogonality). In appendix C we derive the equalities for the type 2 Tobit model and similar logic can be applied to other two-part models. Additional testable results for the aforementioned models are discussed in Appendix D.

### v2

@@ Abstract.
This paper develops sharp testable implications for Tobit and IV-Tobit models’ identifying assumptions: linear index specification, (joint) normality of latent errors, and treatment (instrument) exogeneity and relevance. The new sharp testable equalities can detect all possible observable violations of the identifying conditions. We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds. Simulation results suggest adequate test size for large samples and reasonable power in detecting violations of the exogeneity assumption and error structure. We review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions. We revisit a study of married women’s labor supply in Lee (1995) to demonstrate the practical implementation and usefulness of the test.
The present version is as of January 17, 2025, the first version is from December 2022. All errors are ours. Email address: otavio.bartalotti@monash.edu. We want to thank participants at seminars in the 2024 Jornadas Anuales de Economía del Banco Central del Uruguay and RMIT University, the 2024 IAAE and Southern Economic Association meetings.
@@ 1. Introduction
Since the seminal work of Tobin (1958), Tobit models have earned attention in economics, business and social sciences. [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations. Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative. This approach is related to a broader class of censored or truncated regression models, depending on whether observations outside a specified range are lost or censored. When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE). Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument. While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
@@
In this paper, we develop a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature. The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model. Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability. Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013). The Tobit family of models involves continuous outcomes with accumulation points and potentially continuous treatment, implying a large number of moment equalities to be tested. This creates significant challenges to test implementation. We propose a discretization of the space of the treatment and the outcome that balances the computational requirements and data availability for different parts of their joint support. This simplifies the implementation, making it easy to compute and providing an asymptotically valid testing procedure.
The third contribution is to review and propose alternative approaches that can be used when the model is rejected. We explore an alternative path to partially identify the parameter of interest by assuming the monotonicity of selection into treatment. Finally, we provide an empirical example illustrating the methodology’s practical relevance. More generally, the current paper contributes to the growing literature on testing identifying assumptions of econometric models.
@@
We focus on two main models: (i) the “classic Tobit” model in which the main variable of interest is assumed to be exogenous and (ii) the instrumental variable (IV) Tobit model. In both cases, the proposed test considers all observable violations of the model structure, such as linear index, normality of the latent errors, independence of the treatment, and homoskedasticity; with the addition of the validity of the instrument for the IV-Tobit. Additional results for variants of the Tobit family of models are presented in the Appendix.
@@ 1.1. Previous Literature
There is a vast literature related to testing the validity of Tobit models and their assumptions, which we contribute to. Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid. Nelson (1981) constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates. Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model. Bera, Jarque, and Lee (1984) state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained. Newey (1987b) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test. Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the LM (score) test for a common null hypothesis. Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
@@
Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model. With a similar intuition to our framework, their test writes down conditional moment restrictions, which should have zero conditional expected values under the null. Since the model was estimated by maximum likelihood, the assumed data-generating process specifies the moments of disturbances conditional on the covariates to be the ones of a normal distribution. Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution. While intuitively similar, our procedure detects all possible violations of the model, not only those evident from deviations in the third and fourth moments. Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables. Most of these approaches focus on testing a particular subset of assumptions or a specific class of alternatives to consider.
@@
The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data. This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models. Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, encouragement designs, among others. (Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023; Bai and Tabord-Meehan, 2024). Our testing procedures are connected to Kédagni and Mourifié (2020), which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
The two papers closest to ours are Acerenza, Bartalotti, and Kédagni (2023) and Goff, Kédagni, and Wu (2024). The first focuses on bivariate probit models, leading to a finite set of moment equalities as testable restrictions. As mentioned above, this study builds upon their work by tackling the case of the Tobit family of estimators, which provides additional technical and practical challenges given the continuous nature of outcome and treatment, including accumulation points. Goff, Kédagni, and Wu (2024) considers separable parametric instrumental variable models that induce a different set of moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable. [^2] While Goff, Kédagni, and Wu (2024) briefly mention the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness. Implementation would also differ from our approach.
@@
The remainder of the paper is organized as follows. Section 2 presents the model and identifying assumptions. Section 3 derives the sharp testable implications. Section 4 discusses heuristically the identification of the model’s parameters. Section 5 outlines the testing procedure. Section 6 include simulation evidence about the test’s size while discussions about power are relegated to the Appendix. Section 7 discusses how to relax the assumptions in case of rejection. Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from Lee (1995) to demonstrate the practical implementation and usefulness of the test. Finally, Section 9 concludes. Additional results and details are collected in Appendices A-F.
@@ 2. Models
This section describes two popular “two-part” models for truncated data for which we derive testable implications in Section 3: the classic Tobit, and the IV-Tobit. [^3] We omit conditioning on other exogenous covariates ($X$) for simplicity of exposition. The results including covariates are presented in Section 5.
@@ 2.1. Classic Tobit model
where $U$ is an unobservable (latent) error. In addition to the model structure in system 2.4, the classic Tobit restricts the distribution of the error term and its relationship to the treatment.
@@
Assumption 1 states that the treatment of interest is independent of the model’s unobservables. Assumption 2 imposes that the latent error has a standard normal distribution. [^5] The error could take $N(0,\sigma^{2})$ but after scale normalizations we can model it as $N(0,1)$.
Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatment, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
@@ 2.2. IV Tobit model
The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect. Let the IV Tobit model be: [^6] Results for the IV-Tobit model are presented for continuous $D$ but hold for a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$. In that case, the parameters are identified only up to scale.
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.8)
@@
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.12)
Where $U,V$ are the latent structural error terms. To that model structure, restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable are added.
@@
$Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
@@
Assumption 4 characterizes the distribution of the latent vector of error terms as bivariate normal. Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. [^7] Note that if $U,V$ follow a bivariate normal distribution, $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution, $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$. Then the reduced form system can be scale normalized. The latter characterization aids the development of the sharp testable implications.
@@ 3. Sharp testable equalities
This section presents the sharp testable equalities for the models described in Section 2. Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, independence, instrument validity, and normality) is valid. The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 4.
@@ 3.1. Classic Tobit model
Starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ are observed for $c_{1},c_{0}>0$. For any value of the treatment variables $d\in\mathcal{D}$:
@@
Turning to the accumulation point, the observed event of $Y=0$ has probability,
$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi(-\alpha_{0}-\alpha_{1}d).$ (3.3)
@@
In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$. For the proof, see Appendix A.
@@
The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable. However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities 3.1-3.3. See Appendix E for the proof.
@@
The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (Barros et al., 2018), different thresholds (Carson and Sun, 2007), dynamic Tobit models, or including individual specific effects (Wooldridge, 2005; Honore, Kyriazidou, and Powell, 2000; Honoré, 1993). In appendix C, we derive the equalities for the type 2 Tobit model, and similar logic can be applied to other two-part models. Additional testable results for the aforementioned models are discussed in Appendix D.